A student solved the quadratic equation by finding the square root of both sides. In which line is a mistake first made?
b^2 - 4b + 4 = 16
Line 1:(b - 2)^2 = 16
Line 2:b - 2 = 16
Line 3:b = 18

Answer by Alan3354(69443)

A student solved the quadratic equation by finding the square root of both sides. In which line is a mistake first made?
b^2 - 4b + 4 = 16
Line 1:(b - 2)^2 = 16
Line 2:b - 2 = 16 *********** should be sqrt(16) = ±4
Line 3:b = 18